# SparseRadNet: Sparse Perception Neural Network on Subsampled Radar Data

### v2

@@ 3 SparseRadNet architecture
Figure 2 shows the architecture of our SparseRadNet, which has four modules ordered from the network’s input to its output: (i) the deep radar subsampling module for selecting important parts from the input RD spectrum (Sec. 3.1); (ii) the two-branch backbone with an attentive fusion module for capturing neighbor information (Sec. 3.2); (iii) the range-angle decoder for transforming from RD to RA view (Sec. 3.3); and (iv) two output heads for object detection and freespace segmentation (Sec. 3.4).
@@ 3.1 Deep Radar Subsampling Module
where $\text{detach}(\cdot)$ removes a tensor from the computational graph. In the forward pass, mask $\mathbb{A}$ serves as a binary sampling mask and produces $\mathbb{y}$ as the sparse input to downstream perception modules (black solid arrow in Fig. 2a). While in the backward pass, the gradient propagates to our CNN layers through the second term in Eq. 3 (red dashed arrow in Fig. 2a).
@@ 4.1 Implementation details
In the DRS module, we configure the number of sampled pixels as $M=4000$, which constitutes only 3% of the input. To reduce the memory required during training and lighten the burden on the softmax, we perform average pooling and sample 1000 $2\times 2$ patches. The temperature parameter $\tau$ of the softmax is set to 4. To speed up the inference, we do not add Gumbel noise in the evaluation mode and only compute the hard mask $\mathbb{A}^{hard}$ for a forward pass. We observe that a sparse backbone cannot provide full gradient information for the softmax function when gradients backpropagate to $\mathbb{A}^{soft}$, while a dense backbone enables the training of the DRS module. Meanwhile, the MIMO pre-encoder could incorporate prior knowledge about the repeated object signatures into the gradients. Therefore, we pre-train our DRS module by plugging it into the dense baseline model FFT-RadNet [34]. Then we load and freeze the weights of this module when training our SparseRadNet.
@@
We use the Adam optimizer with a learning rate of $10^{-4}$ and a decay weight of 0.9 for every 10 epochs. We train for 100 epochs and pick the model that achieves the highest average F1 score on the validation set. More details about the baseline model, loss functions, evaluation metrics, and detailed ablation study set-up can be found in the supplement.

### v3

@@ 3 SparseRadNet architecture
Figure 2 shows the architecture of our SparseRadNet, which has four modules: (i) the deep radar subsampling module for selecting important parts from the input RD spectrum (Sec. 3.1); (ii) the two-branch backbone with an attentive fusion module for capturing neighbor information (Sec. 3.2); (iii) the range-angle decoder for transforming from RD to RA view (Sec. 3.3); and (iv) two output heads for object detection and freespace segmentation (Sec. 3.4).
@@ 3.1 Deep Radar Subsampling Module
where $\text{detach}(\cdot)$ removes a tensor from the computational graph. In the forward pass, mask $\mathbb{A}$ serves as a binary sampling mask and produces $\mathbb{y}$ as the sparse input to downstream perception modules (black solid arrow in Fig. 2a). While in the backward pass, gradients propagates to Eq. 3 where only the second term $\mathbb{A}^{soft}$ requires gradient. In Eq. 2, gradients are distributed by the differentiable proxy $\text{softmax}_{\tau}$ to all logits $\mathbb{Z}_{\theta}(\mathbb{x})$, then to our sampling layers’ parameters $\theta$ (red dashed arrow in Fig. 2a).
@@ 4.1 Implementation details
In the DRS module, we configure the number of sampled pixels as $M=4000$, which constitutes only 3% of the input. To reduce the memory required during training and lighten the burden on the softmax, we perform average pooling and sample 1000 $2\times 2$ patches. The temperature parameter $\tau$ of the softmax is set to 4. To speed up the inference, we do not add Gumbel noise in the evaluation mode and only compute the hard mask $\mathbb{A}^{hard}$ for a forward pass. A sparse backbone cannot provide complete gradient information for the $\text{softmax}_{\tau}$ over all pixels when gradients backpropagate to $\mathbb{A}^{soft}$, while a dense backbone enables the training of the DRS module. Meanwhile, the MIMO pre-encoder could incorporate prior knowledge about the repeated object signatures into the gradients. Therefore, we pre-train our DRS module by plugging it into the dense baseline model FFT-RadNet [34]. Then we load and freeze the weights of this module when training our SparseRadNet.
@@
We use the Adam optimizer with a learning rate of $10^{-4}$ and a decay weight of 0.9 for every 10 epochs. We train for 100 epochs and pick the model that achieves the highest average F1 score on the validation set. More details about the baseline model, loss functions, evaluation metrics and detailed ablation study set-up can be found in the supplement.
@@ 5 Conclusion
Acknowledgements
J.W. and M.R. acknowledge support by the German Federal Ministry of Education and Research within the junior research group project “UnrEAL” (grant no. 01IS22069).